Find the negation of each: 1.



x ( (x+3=2) ^ (x


-


2 =1))







1
Expert's answer
2021-01-25T02:20:37-0500

"\nexists x ( (x+3=2) \land (x-2 =1)) = \forall x (\lnot(x+3=2) \lor \lnot (x-2 =1)) = \forall x ((x+3\ne2) \lor (x-2 \ne 1)) = \forall x (x\ne -1\lor x\ne 3)"